OpenStudy (anonymous):

Use and write different forms of linear equations, such as slope-intercept, point-slope, or standard form. Be sure you are able to identify which form is best in specific situations.

OpenStudy (anonymous):

someone please help

OpenStudy (anonymous):

i will reward medals.

jigglypuff314 (jigglypuff314):

slope-intercept form: y = mx + b when m = slope and b = y-intercept point slope form: y - y1 = m(x - x1) when m = slope and given the point (x1, y1) standard form: ax + by = c when a b and c are integers

OpenStudy (anonymous):

so is that a forumula?

OpenStudy (anonymous):

oh i have to have a situation for one of them.

jigglypuff314 (jigglypuff314):

those are the formulas for which you would plug into :) so for example if given slope is 2 and y-intercept is (0, 3) then you can use slope-intercept form: y = mx + b plug in slope = m = 2 and b = y-intercept = 3 so for that example the equation would be y = 2x + 3

OpenStudy (anonymous):

uh i need a real life situation with this.

OpenStudy (anonymous):

jigglypuff?

jigglypuff314 (jigglypuff314):

I'm not too good at thinking up this sort of stuff ;) number of hours past: 0, 1, 2, 3 number of files copied: 5, 8, 11, 14 for this situation, every 1 hour, 3 copies are made and when it was 0 hours, there had been 5 copies therefore m = 3 and b = 5 so equation would be y = 3x + 5

OpenStudy (anonymous):

ok that works thank you so much! :-)

jigglypuff314 (jigglypuff314):

glad I could help :)